pls circle correct answer
Determine the value of Voltage V2. 0.6 A 12 v 1322 + 112 3 12 0.125 V 0.525 V 1.125 V 1.925 V 2.925 V
Determine the value of the Current 13. 12 V + - 9V + W - 200 1 3622 3 \13 0.15 A 0.45 A 1.05 A 0.44 A 0.35 A
Answer #1

Given circuit 62 ^ZI The A To find a 12-12 = 104 ) 12-V2=8.4 12=3.60 - 0.6A - 132 t ANASA 12v {ur V2 Applying Source transforVq=4.2011 24 112=1925v ) + lov a w 362 202 & I3 ju for V, apply kve for outer loop -0, +12+9 = 0 0, -210 + 124_ V3 + qv - 200V₂=9v Applying Nodal Analysis at v₃ 2001 + Iz = 0.6-0.25 Iz = 0.35 A